The train traveled 300 kilometers north to Chicago in 3 hours. What is the

train's speed? What is the train's velocity?*


1
Expert's answer
2021-02-12T18:33:04-0500

The speed is a scalar quantity and it has the magnitude only. The speed of the train can be found as follows:


"v=\dfrac{s}{t}=\dfrac{300\ km}{3\ h}=100\ \dfrac{km}{h}."

The velocity is a vector quantity and it has both magnitude and direction. Therefore, the train's velocity is "100\ \dfrac{km}{h} [North]".